OpenStudy (anonymous):

Find all rational roots of 2x^3 - 7x^2 + 9 = 0.

Nnesha (nnesha):

do you know p/q method ?

OpenStudy (anonymous):

@Nnesha - no, sorry.

Nnesha (nnesha):

\[\huge\rm \frac{ p }{ q } =\frac{ constant }{ leading ~cofficient }\] what is constant and leading coefficient in that equation ??

OpenStudy (anonymous):

the constant term is 9 and the leading cofficient is 2.

Nnesha (nnesha):

yes right so p represent constant q = leading coefficient so you ahve \[\frac{ 9 }{ 2 } = \frac{ \pm 1, ??? }{ ? }\] now factor both 9 and 2 write with plus minus sign for example 1

OpenStudy (anonymous):

9 = 1, 2 ,3 2 = 1, 2

Nnesha (nnesha):

yep and 9 too

Nnesha (nnesha):

so you ahve \[\huge\rm \frac{ 9 }{ 2 } = \frac{ \pm 1, \pm 3 , \pm 9??? }{ \pm 1 , \pm 2}\] 2 is not a factor of 9

Nnesha (nnesha):

now write everything in fraction form  for example 1 divide by both 1 nd 2 which is at the denominator \[\huge\rm \frac{ \pm 1 }{ \pm 1 } ,\frac{ \pm 1 }{ \pm 2}\]